OpenStudy (anonymous):

The equation x2+2ix-4=0 has its roots

OpenStudy (anonymous):

x=sqrt(3)-i and x=-sqrt(3)-i

OpenStudy (anonymous):

x=sqrt(3)-i, ≈1.7320508075689-i x=-sqrt(3)-i, ≈-1.7320508075689-i